# Efficient Prior Selection in Gaussian Process Bandits with Thompson Sampling

### v1

@@ 3 Algorithms
As discussed by Russo & Roy (2014), TS can offer advantages over UCB algorithms for problems where constructing tight confidence bounds is difficult. In addition, Thompson sampling is often observed to perform better than UCB in practice Chapelle & Li (2011); Wen et al. (2015); Kandasamy et al. (2018); Åkerblom et al. (2023b, a). Motivated by this, we present two algorithms for efficient prior selection based on TS.
@@ 3.1 Prior-Elimination with Thompson sampling
Figure 2: Elimination procedure of PE-GP-TS. The solid lines correspond to posterior means and the shaded regions are confidence intervals. The dashed lines are samples from the posteriors. The figure has been adapted from Ziomek et al. (2024b) to illustrate the elimination step of PE-GP-TS.
@@
3:     Sample $\tilde{f}_{t,p}\sim{\mathcal{GP}}(\mu_{t,p},k_{t,p})$ $\forall p\in P_{t-1}$
@@
12:        $P_{t}=P_{t-1}$
Our first algorithm is an extension of the PE-GP-UCB algorithm Ziomek et al. (2024b) to be employed with Thompson sampling - instead of UCB. The key difference is that instead of maximizing the upper confidence bound $U_{t}(x,p)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$ over ${\mathcal{X}}\times P$, we instead sample $\tilde{f}_{t,p}$ from the posterior ${\mathcal{GP}}(\mu_{t,p},k_{t,p})$ for all priors $p\in P_{t-1}$ where $P_{t-1}$ is the set of active priors. Then, we select the arm and prior $x_{t},p_{t}$ such that $x_{t},p_{t}=\operatorname*{arg\,max}_{x,p\in{\mathcal{X}}\times P_{t-1}}\tilde{f}_{t,p}(x)$. Whilst PE-GP-UCB has two layers of optimism, the upper confidence bound and joint maximization of $x$ and $p$. PE-GP-TS has only a single layer of optimism - which should alleviate potential overexploration issues.
The elimination procedure of PE-GP-TS is illustrated in Fig. 2. Samples $\tilde{f}_{t,p}$ are drawn from the active prior $p\in P_{t-1}$. Then, the unknown function $f$ is queried at the selected arm $x_{t}$. If the observed value differs too much from the prediction made by the selected prior, then the selected prior is eliminated. Otherwise, it remains active.
The PE-GP-TS algorithm is presented in Algorithm 1. Similar to PE-GP-UCB, the set $S_{t,p}$ is used to store the time steps where prior $p$ was selected up to and including time $t$. When prior $p_{t}$ is selected, the prediction error $\eta_{t}=y_{t}-\mu_{t,p_{t}}(x_{t})$ between the observed and predicted value made by the prior $p_{t}$ is computed. If the sum of prediction errors made by the prior $p_{t}$ exceeds the threshold value $V_{t}$, then $p_{t}$ is eliminated from the active priors $P_{t}$, see line 9. Note that at time step $t$, only the selected prior $p_{t}$ can be eliminated. As such, if a prior is very pessimistic it may never be selected and therefore will never be eliminated. Thus, the final set of active priors $P_{T}$ should be viewed as non-eliminated priors rather than necessarily being reasonable priors.
@@ 3.2 HyperPrior Thompson sampling
0:  Horizon $T$, prior functions $\{\mu_{1,p},k_{1,p}\}_{p\in P}$, hyperprior $P_{0}$.
@@
6:     Set ${P}_{t}(p)\propto{\mathbb{P}}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)P_{t-1}(p)$triangleright Update hyperposterior
@@
The algorithm is presented in Algorithm 2. In the first step, the current prior $p_{t}$ is sampled from the hyperposterior $P_{t-1}$. Then, a single sample $\tilde{f}_{t}$ is taken from the selected posterior ${\mathcal{GP}}(\mu_{t,p_{t}},k_{t,p_{t}})$ and is used to select the current arm: $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}\tilde{f}_{t}(x)$. After observing $y_{t}$, the hyperposterior is updated by computing the likelihood of $y_{t}$ under the different priors. Note that since the set of priors $P$ is finite, computing the posterior is tractable albeit computationally costly with a complexity of ${\mathcal{O}}(t^{3}|P|)$. The likelihood ${\mathbb{P}}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)={\mathcal{N}}(y_{t};\mu_{t,p}(x_{t}),\sigma_{t,p}^{2}(x_{t})+\sigma^{2})$ is simply the Gaussian likelihood of the posterior at $x_{t}$ plus Gaussian noise with variance $\sigma^{2}$.
4 Regret bounds
In this section, we present regret bounds for the proposed algorithms. Recall from the problem statement that we consider two slightly different settings for the two algorithms. Specifically, for PE-GP-TS we assume the unknown prior $p^{*}$ is sampled arbitrarily from $P$ whilst for HP-GP-TS we assume that the unknown prior $p^{*}$ is selected from a known hyperprior distribution $P_{0}$.
Ziomek et al. (2024b) structured the proof of the regret bound of PE-GP-UCB into 4 larger steps; First, showing that $p^{*}$ is never eliminated with high probability. Second, establishing a bound on the simple regret. Third, bounding the cumulative regret. Finally, the cumulative bound is re-expressed in terms of the worst-case MIG. For PE-GP-TS, we note that it is sufficient to establish a new bound on the simple regret and then adapt the steps of Ziomek et al. to accommodate the new simple regret bound.
@@ 4 Regret bounds
[] For any $p^{*}\in P$, if $f(x)\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ and $\beta_{t}=2\log\left(\frac{|{\mathcal{X}}||P|\pi^{2}t^{2}}{3\delta}\right)$. Then, with probability at least $1-\delta$, the following holds for all $t,x\in[T]\times{\mathcal{X}}$:
@@
All proofs can be found in Appendix A. Section 4 is based on Lemma 5.1 of Srinivas et al. (2012) but adapted to TS by specifying that it holds for any sequence of $x_{1},\ldots,x_{T}$, as discussed by Russo & Roy (2014). Additionally, we add Eq. 5 which can be shown through the same steps and an additional union bound over $P$. Next, we state our bound for the simple regret of PE-GP-TS.
@@
[] If the event of Section 4 holds, then the following holds for the simple regret of PE-GP-TS for all $t\in[T]$:
@@
Compared to the simple regret bound for PE-GP-UCB, we obtain the additional term $2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})$. Since $p^{*}$ is fixed, the sum over $t$ of the new term is ${\mathcal{O}}(\sqrt{T\beta_{T}\gamma_{T,p^{*}}})$ and we obtain the following regret bound:
@@
If $p^{*}\in P$ and $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, then PE-GP-TS with confidence parameters $\beta_{t}=2\log(2|{\mathcal{X}}||P|\pi^{2}t^{2}/3\delta)$ and $\xi_{t}=2\sigma^{2}\log(|P|\pi^{2}t^{2}/3\delta)$, satisfies the following regret bound with probability at least $1-\delta$:
$R(T)\leq$ $2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}$ (7)
$+2\sqrt{CT\beta_{T}\gamma_{T,p^{*}}}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}$
where $B_{p^{*}}=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$ and $C=2/\log(1+\sigma^{-2})$.
Since $\gamma_{T,p^{*}}\leq\hat{\gamma}_{T}$, the bound is of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma}_{T}})$ w.r.t. $T$ which matches that of PE-GP-UCB. To our knowledge, the best lower bound for standard GP bandits in the Bayesian setting, where $f$ is sampled from a GP, is $\Omega(\sqrt{T})$ for $d=1$ Scarlett (2018). This would suggest that our bound is tight up to a factor ${\mathcal{O}}(\sqrt{\beta_{T}\hat{\gamma}_{T}})$.
[…]
{lemma}
[] Let $C=2/\log(1+\sigma^{-2})$ and assume $\sigma_{t,p}(x)\leq 1$, then for any sequence of arms $x_{1},\ldots,x_{T}\in{\mathcal{X}}$ and any $p\in P$
$I_{p}({\mathbf{y}}_{T};f)=\frac{1}{2}\sum_{t\in[T]}\log\left(1+\sigma^{-2}\sigma_{t,p}^{2}(x_{t})\right),$ (8)
${\sum_{t\in[T]}}\sigma^{2}_{t,p}(x_{t})\leq CI_{p}({\mathbf{y}}_{t};f).$ (9)
{corollary}
[] For any stochastic or deterministic sequence of arms $x_{1},\ldots,x_{T}\in{\mathcal{X}}$ and $p^{*}\sim P_{0}$,
${\mathbb{E}}\Big{[}{\sum_{t\in[T]}}\sigma_{t,p^{*}}^{2}(x_{t})\Big{]}\leq C\bar{\gamma_{T}}.$ (10)
The proof of Theorem 4.1 is based on Lemma 5.3 and 5.4 in Srinivas et al. (2012) but is adapted to randomized selection by specifying that it holds for any sequence of arms. Theorem 4.1 follows by Theorem 4.1 and the definition of maximum information gain. Next, we state our regret bound for HP-GP-TS.
@@
If $p^{*}\sim P_{0}$ and $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, then the Bayesian regret of HP-GP-TS is bounded by
${\text{BR}}(T)\leq\frac{\pi^{2}}{6}+\sqrt{C_{1}T\beta_{T}\bar{\gamma}_{T}}$ (11)
where $C_{1}=2\log(1+\sigma^{-2})$ and $\beta_{t}=2\log(|{\mathcal{X}}|t^{2}/\sqrt{2\pi)}$.
Unlike PE-GP-TS and PE-GP-UCB, the regret bound for FB-GP-TS depends on the average $\bar{\gamma}_{T}$ rather than the worst case $\hat{\gamma}_{T}$ which can impact the theoretical regret significantly if the complexity of learning the priors differ. Although, this is reasonable since the elimination methods assume arbitrary selection of $p^{*}$ as opposed to sampling from a hyperprior. If the hyperprior is deterministic then the regret bound for FB-GP-TS matches that of GP-TS up to a factor ${\mathcal{O}}(\sqrt{\log T})$ Takeno et al. (2024) and the average $\bar{\gamma}_{T}$ would be equal to the worst case $\hat{\gamma}_{T}$. Again, using the lower bound of Scarlett (2018), our upper bound would tight up to a factor of ${\mathcal{O}}(\sqrt{\beta_{T}\bar{\gamma}_{T}})$.
Figure 3: Prior mean functions for the hill experiment. Mean function 0 consists of 10 smaller hills whereas mean functions 1 to 4 have 9 smaller hills and 1 taller hills.
Figure 4: Cumulative regret for synthetic experiments with varying kernel (left), lengthscale (center) and mean function (right). The average final regret for PE-GP-UCB is 116.5 in the lengthscale experiment. Errorbars correspond to ±$\pm 1$ standard error.
Figure 5: Average amount of priors remaining in $P_{t}$ over time for PE-GP-UCB and -TS.

### v2

@@ 3 Algorithms
As discussed by Russo & Van Roy (2014), TS can offer advantages over UCB algorithms for problems where constructing tight confidence bounds is difficult. In addition, Thompson sampling is often observed to perform better than UCB in practice (Chapelle & Li, 2011; Wen et al., 2015; Kandasamy et al., 2018; Åkerblom et al., 2023b, a). Motivated by this, we present two algorithms for efficient prior selection based on TS.
@@ 3.1 Prior-Elimination with Thompson sampling
3:  Sample $\tilde{f}_{t,p}\sim{\mathcal{GP}}(\mu_{t,p},k_{t,p})$ $\forall p\in P_{t}$
@@
12:   $P_{t+1}=P_{t}$
Our first algorithm is an extension of PE-GP-UCB (Ziomek et al., 2025) to be employed with Thompson sampling - instead of UCB. The key difference is that instead of maximizing the upper confidence bound $U_{t}(x,p)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$ over ${\mathcal{X}}\times P_{t-1}$, we instead sample $\tilde{f}_{t,p}$ from the posterior ${\mathcal{GP}}(\mu_{t,p},k_{t,p})$ for all priors $p\in P_{t-1}$ where $P_{t-1}$ is the set of active priors. Then, we select the arm and prior $x_{t},p_{t}$ such that $x_{t},p_{t}=\operatorname*{arg\,max}_{x,p\in{\mathcal{X}}\times P_{t-1}}\tilde{f}_{t,p}(x)$. Whilst PE-GP-UCB has two layers of optimism, the upper confidence bound and joint maximization of $x$ and $p$, PE-GP-TS has only a single layer of optimism - which should alleviate potential overexploration issues.
The elimination procedure of PE-GP-TS is illustrated in Fig.˜2. Samples $\tilde{f}_{t,p}$ are drawn from the active prior $p\in P_{t-1}$. Then, the unknown function $f$ is queried at the selected arm $x_{t}$. If the observed value differs too much from the prediction made by the selected prior, then the selected prior is eliminated. Otherwise, it remains active.
The PE-GP-TS algorithm is presented in Algorithm˜1. Similar to PE-GP-UCB, the set $S_{t,p}$ is used to store the time steps where prior $p$ was selected up to and including time $t$. When prior $p_{t}$ is selected, the prediction error $\eta_{t}=y_{t}-\mu_{t,p_{t}}(x_{t})$ between the observed and predicted value made by the prior $p_{t}$ is computed. If the sum of prediction errors made by the prior $p_{t}$ exceeds the threshold value $V_{t}$, then $p_{t}$ is eliminated from the active priors $P_{t}$, see line˜9. Note that at time step $t$, only the selected prior $p_{t}$ can be eliminated. As such, if a prior is very pessimistic it may never be selected and therefore will never be eliminated. Thus, the final set of active priors $P_{T}$ should be viewed as non-eliminated priors rather than necessarily being reasonable priors.
@@ 3.2 HyperPrior Thompson sampling
0: Horizon $T$, prior functions $\{\mu_{1,p},k_{1,p}\}_{p\in P}$, hyperprior $P_{1}$.
@@
6:  Set ${P}_{t+1}(p)\propto\mathbb{P}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)\cdot P_{t}(p)$ $\triangleright$ Update hyperposterior
@@
The algorithm is visualized in Fig.˜2 and presented with more details in Algorithm˜2. In the first step, the current prior $p_{t}$ is sampled from the hyperposterior $P_{t-1}$. Then, a single sample $\tilde{f}_{t}$ is taken from the selected posterior ${\mathcal{GP}}(\mu_{t,p_{t}},k_{t,p_{t}})$ and is used to select the current arm: $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}}\tilde{f}_{t}(x)$. After observing $y_{t}$, the hyperposterior is updated by computing the likelihood of $y_{t}$ under the different priors. Note that since the set of priors $P$ is finite, computing the posterior is tractable albeit computationally costly with a complexity of ${\mathcal{O}}(t^{3}|P|)$. The likelihood $\mathbb{P}(y_{t}|x_{t},\{x_{i},y_{i}\}_{i=1}^{t-1},p)={\mathcal{N}}(y_{t};\mu_{t,p}(x_{t}),\sigma_{t,p}^{2}(x_{t})+\sigma^{2})$ is simply the Gaussian likelihood of the posterior at $x_{t}$ with added Gaussian noise with variance $\sigma^{2}$.
4 Regret analysis
In this section, we analyze the regret for the proposed algorithms. Recall from the problem statement that we consider two slightly different settings for the two algorithms. Specifically, for PE-GP-TS we assume the unknown prior $p^{*}$ is selected arbitrarily from $P$ whilst for HP-GP-TS we assume that the unknown prior $p^{*}$ is selected from a known hyperprior distribution $P_{1}$.
4.1 Analysis of PE-GP-TS
Ziomek et al. (2025) structured the proof of the regret bound of PE-GP-UCB into 4 larger steps; First, showing that $p^{*}$ is never eliminated with high probability. Second, establishing a bound on the simple regret. Third, bounding the cumulative regret. Finally, the cumulative bound is re-expressed in terms of the worst-case MIG. For PE-GP-TS, we establish a new bound on the simple regret and then adapt the steps of Ziomek et al. to accommodate the new simple regret bound.
@@ 4.1 Analysis of PE-GP-TS
[] If $f(x)\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ and $\beta_{t}=2\log\left(\frac{|{\mathcal{X}}||P|\pi^{2}t^{2}}{3\delta}\right)$. Then, with probability at least $1-\delta$, the following holds for all $t,x,p\in[T]\times{\mathcal{X}}\times P$:
@@
All proofs can be found in Appendix˜A. Section˜4.1 is based on Lemma 5.1 of Srinivas et al. (2012) but adapted to TS by specifying that it holds for any sequence of $x_{1},\ldots,x_{T}$, as discussed by Russo & Van Roy (2014). Additionally, we add Eq.˜5 which can be shown through the same steps and an additional union bound over $P$. Next, we state our bound for the simple regret of PE-GP-TS.
@@
[] If the event of Section˜4.1 holds, then the following holds for the simple regret of PE-GP-TS for all $t\in[T]$:
@@
Compared to the simple regret bound for PE-GP-UCB, we obtain the additional term $2\sqrt{\beta_{t}}\sigma_{t,p^{*}}(x^{*})$ which leads to the following regret bound:
@@
Let $B_{p^{*}}=\beta_{1}+\sup_{x\in{\mathcal{X}}}|\mu_{1,p^{*}}(x)|$ and $C=2/\log(1+\sigma^{-2})$. If $p^{*}\in P$ and $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, then PE-GP-TS with confidence parameters $\beta_{t}=2\log(2|{\mathcal{X}}||P|\pi^{2}t^{2}/3\delta)$ and $\xi_{t}=2\sigma^{2}\log(|P|\pi^{2}t^{2}/3\delta)$, satisfies the following regret bound with probability at least $1-\delta$:
$R(T)\leq 2|P|B_{p^{*}}+2\sqrt{\xi_{T}|P|T}+2\sqrt{CT\beta_{T}\hat{\gamma}_{T}|P|}+2\sqrt{CT\beta_{T}{\sum_{t\in[T]}}\sigma^{2}_{t,p^{*}}(x^{*})}$ (7)
The bound of the first three terms is of order ${\mathcal{O}}(\sqrt{T\beta_{T}\hat{\gamma}_{T}})$ w.r.t. $T$ which matches that of PE-GP-UCB. To our knowledge, the best lower bound for standard GP bandits in the Bayesian setting, where $f$ is sampled from a GP, is $\Omega(\sqrt{T})$ for $d=1$ (Scarlett, 2018). This would suggest that our bound is tight up to a factor ${\mathcal{O}}(\sqrt{\beta_{T}\hat{\gamma}_{T}})$.
4.2 Analysis of HP-GP-TS
We analyze the regret of HP-GP-TS in two ways: using the UCB-based framework of Russo & Van Roy (2014) and the information-theoretic framework of Russo & Van Roy (2016). First, note that HP-GP-TS inherits the probability matching property of GP-TS that $x_{t}|H_{t}\overset{d}{=}x^{*}|H_{t}$ where $\overset{d}{=}$ denotes equal in distribution. In addition, $p_{t}|H_{t}\overset{d}{=}p^{*}|H_{t}$ since $p_{t}$ is sampled from the posterior distribution of $p^{*}$. In the UCB-based framework, we decompose the regret into three terms:
${\mathbb{E}}[f(x^{*})-f(x_{t})]={\mathbb{E}}\Big[\underbrace{f(x^{*})-U_{t,p^{*}}(x^{*})}_{(1)}+\underbrace{U_{t,p^{*}}(x^{*})-U_{t,p^{*}}(x_{t})}_{(2)}+\underbrace{U_{t,p^{*}}(x_{t})-f(x_{t})}_{(3)}\Big]$ (8)
where $U_{t,p}(x)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$. Summing over $t\in[T]$, the first term can be bounded by a constant whilst the third term is bounded by $\sqrt{CT\beta_{T}\bar{\gamma}_{T}}$. Together, these two terms match the Bayesian regret bound for GP-TS with known prior. For the second term, one can utilize that $p^{*},x_{t}|H_{t}\overset{d}{=}p_{t},x^{*}|H_{t}$ to re-express it as ${\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$. Hence, the second term can be seen as the cost of learning the true prior. Intuitively, if the priors are similar then the upper-confidence bounds of the true and selected prior will not differ significantly. Similarly, if the priors are sufficiently distinguishable, one would expect $P_{t}(p^{*})$ to increase quickly from a few samples. In the following lemma, we demonstrate the latter for $|P|=2$ with shared kernel functions. {lemma}[] If $|P|=2$ and the two priors share the same kernel function ($k_{p}=k$ $\forall p\in P$), then for any fixed sequence of arms $x_{1:t}=\{x_{i}\}_{i=1}^{t}$ the posterior probability of the true prior $p^{*}$ satisfies
${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|p^{*}=p,x_{1:t}\right]\geq 1+P_{0}(p)e^{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}\Phi\Big(-\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\Big)-\frac{1}{P_{0}(p)}\Phi\Big(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\Big)$ (9)
where ${\bm{\mu}}\in{\mathbb{R}}^{t}$ such that $({\bm{\mu}})_{i}=\mu_{p^{*}}(x_{i})-\mu_{\tilde{p}}(x_{i})$ for $\tilde{p}\neq p^{*}$ and $({\bm{\Sigma}})_{i,j}=k(x_{i},x_{j})$. From Eq.˜8, we note that $P_{t}(p^{*})$ increases the most when $\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert$ is maximized, which corresponds to selecting arms with low correlation and large difference in prior mean (relative to the variance). Hence, $P_{t}(p^{*})$ increases quickly at first when HP-GP-TS explores but levels out when the algorithm starts exploiting. Next, we state the Bayesian regret bound for HP-GP-TS.
@@ 4.2 Analysis of HP-GP-TS
If $p^{*}\sim P_{1}$, $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$ and $\beta_{t}=2\log(|{\mathcal{X}}|t^{2}/\sqrt{2\pi)}$, then the Bayesian regret of HP-GP-TS is bounded by
${\text{BR}}(T)\leq\pi^{2}/6+{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})\right]+\sqrt{CT\beta_{T}\bar{\gamma}_{T}}.$ (10)
Unlike PE-GP-TS and PE-GP-UCB, the regret bound of the third term for HP-GP-TS depends on the average MIG $\sqrt{\bar{\gamma}_{T}}$ rather than the worst case $\sqrt{|P|\hat{\gamma}_{T}}$ which can impact the theoretical regret significantly if the complexity of the priors differ and the prior is weighted towards simple priors. This is reasonable since the elimination methods assume arbitrary selection of $p^{*}$ as opposed to sampling from a hyperprior. If the hyperprior is deterministic then the regret bound for HP-GP-TS matches that of GP-TS up to a factor ${\mathcal{O}}(\sqrt{\log T})$ (Takeno et al., 2024) and $\bar{\gamma}_{T}$ would be equal to the worst case $\hat{\gamma}_{T}$. Again, using the lower bound of Scarlett (2018), our upper bound would be tight up to a factor of ${\mathcal{O}}(\sqrt{\beta_{T}\bar{\gamma}_{T}})$.
The information-theoretic framework of Russo & Van Roy (2016) can be applied generally if the probability matching property is satisfied and the rewards are subgaussian (Vashishtha & Maillard, 2025).
Theorem 4.3.
If $p^{*}\sim P_{1}$, $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, the Bayesian regret of HP-GP-TS is bounded by
${\text{BR}}(T)\leq\sqrt{2|{\mathcal{X}}|\log(|{\mathcal{X}}|)(\sigma^{2}_{0}+\sigma^{2})T}.$ (11)
The proof of Theorem˜4.3 follows the proof in section D.2 of Russo & Van Roy (2016) with subgaussian noise. For completeness, we provide a proof using the GP-bandit notation in Algorithm˜2. The information-theoretic regret bound is ${\mathcal{O}}(\sqrt{T})$ which improves upon the ${\mathcal{O}}(\sqrt{T\beta_{T}\bar{\gamma}_{T}})$ obtained previously and would match the lower bound of Scarlett (2018) in terms of $T$. However, we also obtain a ${\mathcal{O}}(\sqrt{|{\mathcal{X}}|\log|{\mathcal{X}}|})$ dependency.
Figure 3: Cumulative regret for synthetic experiments with varying kernel (left), lengthscale (center) and mean function (right). The average final regret for PE-GP-UCB is 116.5 and 389.0 in the lengthscale and subspace experiments. Errorbars correspond to $\pm 1$ standard error.
Figure 4: Mean number of priors remaining in $P_{t}$ over time for PE-GP-UCB and -TS (left). Mean entropy in the hyperposterior $P_{t}$ over time for HP- and MAP GP-TS (right). The dashed reference values correspond to entropies of discrete distributions with prob. $q$ on one choice and prob. $\frac{1-q}{|P|-1}$ on the other |P|1|P|-1 choices.